Suppose that a firm is selling a good with a marginal cost of $35. Management estimates...

  1. Suppose that a firm is selling a good with a marginal cost of $35. Management estimates demand elasticity to be –2. What is the appropriate price to set in order to maximize profit?
Answer #1

MC = 35 - 9 Ed -2 Laner 3 Monopoly Iudex of =P-MC = 1 P-35 = 1 e ral 2 (P-35) =P 2P-70 =P ap-P = 20 p =20 profitao maximize p